# Beyond Natural Language: LLMs Leveraging Alternative Formats for Enhanced Reasoning and Communication

### v1

@@ Abstract
Natural language (NL) has long been the predominant format for human cognition and communication, and by extension, has been similarly pivotal in the development and application of Large Language Models (LLMs). Yet, besides NL, LLMs have seen various non-NL formats during pre-training, such as code and logical expression. NL’s status as the optimal format for LLMs, particularly in single-LLM reasoning and multi-agent communication, has not been thoroughly examined. In this work, we challenge the default use of NL by exploring the utility of non-NL formats in these contexts. We show that allowing LLMs to autonomously select the most suitable format before reasoning or communicating leads to a 3.3 to 5.7% improvement in reasoning efficiency for different LLMs, and up to a 72.7% reduction in token usage in multi-agent communication, all while maintaining communicative effectiveness. Our comprehensive analysis further reveals that LLMs can devise a format from limited task instructions and that the devised format is effectively transferable across different LLMs. Intriguingly, the structured communication format decided by LLMs exhibits notable parallels with established agent communication languages, suggesting a natural evolution towards efficient, structured communication in agent communication. Our code is released at https://github.com/thunlp/AutoForm.
@@
[…]
@@ 1 Introduction
Recent research challenges the notion that NL is the ideal intermediate format for LLM reasoning and multi-agent communication. Emerging variants of CoT, such as Program-of-Thought (Chen et al., 2022; Gao et al., 2023) and X-of-Thought (Liu et al., 2023) have explored the use of alternative formats like code and mathematical equations, expanding the LLMs’ reasoning capabilities. However, these approaches often integrate external tools, where the alternative formats primarily serve as a means to facilitate tool execution (e.g., prompting LLM to generate code and use code interpreter execution result as the answer). This introduces complexity in discerning whether the performance improvements are attributable to the format itself or the accompanying tools. Additionally, while the natural ambiguities and emotions inherent in NL may align well with the nuances of human communication, these may not be desired in agent communication, where precision is more important. Nonetheless, current multi-agent research predominantly utilizes NL (Li et al., 2023; Wu et al., 2023; Chen et al., 2023), with limited exploration of other potentially more accurate and efficient communication formats.
@@ 2 Related Work
Advances in Large Language Models (LLMs) have led to the development of autonomous agents like AutoGPT (Significant Gravitas, ) and OpenAI Assistant (OpenAI, 2023a), demonstrating success in diverse tasks (Shinn et al., 2023; Mialon et al., 2023; Zhou et al., 2023; Boiko et al., 2023). Recent research extends this to multi-agent systems for collaborative problem-solving (Du et al., 2023; Osika, 2023; Hong et al., 2023; Qian et al., 2023). CAMEL (Li et al., 2023) explores collaborative problem-solving between two agents through role-playing. ChatEval (Chan et al., 2023) assesses model responses using multi-agent debates. AgentVerse (Chen et al., 2023) introduces a comprehensive framework for multi-agent collaboration, highlighting emergent inter-agent behaviors. However, the alternative formats of multi-agent communication remains underexplored and NL is directly adopted across various research. Pham et al. (2023) explores agent communication with hidden states, but is limited to agents with the same open-source LLM. Our work analyze communication formats for both homogeneous and heterogeneous LLMs.
@@ 4 Experiments
[…]
GPT-3.5 CoT 49.0 22.2 36.7 24.6 59.8 58.7 41.8
  +AutoForm 52.5 38.0 39.7 27.6 64.2 61.4 47.2
Gemini Pro CoT 47.0 43.0 23.5 25.1 54.7 - 38.7
  +AutoForm 48.5 54.0 35.3 28.6 55.5 - 44.4
GPT-4 CoT 64.0 94.2 75.0 39.4 78.7 79.9 71.9
  +AutoForm 69.0 98.6 79.4 42.4 80.3 81.5 75.2
@@ 4.1 Experimental Settings
Single-LLM Reasoning
To comprehensively evaluate the impact of model-chosen format as the thought format in the LLMs’ reasoning performance, we select reasoning benchmarks covering different types of reasoning, including logical reasoning (Logic Grid Puzzle (Srivastava et al., 2022) and Information Essentialy (Srivastava et al., 2022)), mathematical reasoning (AQuA (Ling et al., 2017)), causal reasoning (Minute Mysteries QA (Srivastava et al., 2022)) and symbolic reasoning (Coin Flip (Wei et al., 2022)). In our experiments with the AQuA dataset, we adopted two distinct settings: without tool and with tool. The latter follows Chen et al. (2022) and incorporates code interpreter, albeit with minor modifications (Appendix A). For all these tasks, we require the LLMs to generate the answer in a particular format, and we extract the answer with a written regular expression. The accuracy of each dataset is reported. For more details on the experimental settings, please refer to Appendix A.
@@
Multi-Agent Communication
To measure whether alternative formats can streamline communication, we consider scenarios where two agents with different knowledge or contexts are tasked with answering a question. The answer to the question should be derived from the knowledge of only one of the agents, or both agents’ knowledge collectively, therefore requiring information exchange. The two agents speak in turn to discuss and give the final answer. To create such scenarios, we utilize three existing datasets: Hotpot QA (Yang et al., 2018), Wiki Hop (Welbl et al., 2018) and Narrative QA (Kociský et al., 2018). Hotpot QA and Wiki Hop are two multi-hop QA datasets, which require multiple sentences or paragraphs to deduce the final answer. We randomly assign the text segments provided in the datasets to two different agents. Communication is thus needed to derive the correct answer. Narrative QA requires the model to read the whole book and answer a question. The length of a book often exceeds the context limit of the LLMs. We split the books into nearly equal sizes for the two agents, and ask them to answer the question. For evaluation, we use RougeL as the primary metric. More details are elaborated in Appendix A.
4.2 Experimental Results
Single-LLM Reasoning
The comparative efficacy of the AutoForm approach over the conventional Chain-of-Thought (CoT) methodology in single-LLM reasoning tasks is encapsulated in Table 1. We observe clear performance improvements when employing AutoForm across different datasets, compared to the baseline CoT method.
@@ 4.2 Experimental Results
Multi-Agent Communication
The outcomes of our multi-agent communication experiments, detailed in Table 2, provide valuable insights into the efficiency and effectiveness of utilizing alternative communication formats in collaborative environments. In this experiment, we experiment with different model pairings to explore this robustness more comprehensively. And since the initial speaking agent often sets the tone for the communication format, we vary the speaking order in pairings of heterogeneous models, e.g., GPT-4 + GPT-3.5 in the table indicates GPT-4 serves as the initiator. Due to page limit, we place results where GPT-3.5 initiates the conversation at Appendix C.
@@
5 Analysis
Following the presentation of our experimental results, we would like to delve deeper into understanding the implications and nuances of our findings. The experimental results in the last section have clearly demonstrated that allowing LLMs to autonomously decide on the thought or communication format results in enhanced reasoning and more efficient communication. To further elucidate these outcomes, we engage in a comprehensive analysis addressing several pivotal research questions (RQ). These questions aim to unravel the intricacies of format selection by LLMs and its impact on task performance across different scenarios:
@@ 5 Analysis
  • RQ1: What formats are chosen in single-LLM reasoning? (Section 5.1)
  • RQ2: Can LLMs devise a general format for a task based on some task inputs? (Section 5.2)
  • RQ3: Is the decided format transferable between different LLMs? (Section 5.3)
  • RQ4: What are the features of the formats used in multi-agent communication?  (Section 5.4)
  • RQ5: Does the autonomously determined multi-agent communication format align with conventional agent communication languages such as KQML (Finin et al., 1994)? (Section 5.5)
Model Logic Grid Coin Flip Info Essen MM QA AQuA Average
GPT-3.5 (Instance-Based) 48.0 49.2 39.7 26.1 66.5 45.9 ${}_{(+7.5\%)}$
GPT-3.5 (Task-Based) 51.0 62.8 42.6 28.1 65.0 49.9 ${}_{(+11.5\%)}$
Gemini Pro (Instance-Based) 39.5 44.8 35.3 27.1 59.4 41.2 ${}_{(+2.5\%)}$
Gemini Pro (Task-Based) 41.5 47.8 35.3 28.6 59.4 42.5 ${}_{(+3.8\%)}$
GPT-4 (Instance-Based) 71.5 100.0 76.5 41.4 78.3 73.5 ${}_{(+3.2\%)}$
GPT-4 (Task-Based) 70.0 99.8 75.0 41.4 79.5 73.1 ${}_{(+2.8\%)}$
GPT-4 $\Rightarrow$ GPT-3.5 (Task-Based) 47.5 83.0 35.3 23.2 59.4 49.7 ${}_{(+11.2\%)}$
GPT-4 $\Rightarrow$ Gemini Pro (Task-Based) 40.0 74.6 25.0 25.6 47.2 42.5 ${}_{(+3.8\%)}$
GPT-3.5 $\Rightarrow$ GPT-4 (Task-Based) 65.5 98.6 72.1 42.9 79.5 71.7 ${}_{(+1.3\%)}$
5.1 RQ1: Formats Chosen in Single-LLM Reasoning
@@
5.2 RQ2: Generalization of Format Selection Based on Task Inputs
@@
5.3 RQ3: Transferability of Format Across Different LLMs
@@
The results are presented in the last group of Table 3. Generally, the format is transferable, but may lead to slightly inferior performances compared to the homogeneous setting. For example, transferring the format decided by GPT-4 to GPT-3.5 or Gemini Pro leads to a decrease compared to the homogeneous setting for the two LLMs in most of the tasks except for Coin Flip. On the other hand, when transferring the format decided by GPT-3.5 to GPT-4, the results are generally comparable to the homogeneous setting for GPT-4. For information essentiality, the format selected by GPT-3.5 is generally less efficient, leading to incomplete problem-solving processes. We give examples in Tables 32, 38 and 35. For other tasks, the format generated by GPT-3.5 proves adequate and is easily interpreted by GPT-4, resulting in similar performances to the homogeneous setting for GPT-4.
5.4 RQ4: Features of Communication Format
@@
Brevity and Efficiency. Another prominent feature is the focus on brevity, underlining efficiency in communication. The formats chosen by LLMs often omit elements typically associated with human interaction, such as conventional pleasantries or emotive expressions, resulting in more concise exchanges. This brevity not only conserves computational resources but also concentrates the dialogue on the task at hand. By eliminating extraneous elements, the agents optimize the communication process, leading to faster and more efficient information exchange. This could be beneficial in scenarios where rapid and effective decision-making is paramount.
5.5 RQ5: Alignment with Conventional Agent Communication Languages
@@
Hotpot QA Multi Wiki Narrative QA
@@
The implications of these results are twofold. Firstly, they affirm the potential of AutoForm in generating communication patterns that parallel the formal structures of established ACLs. Secondly, they demonstrate the efficiency of AutoForm in distilling the essence of these traditional formats into a more concise form, thereby conserving computational resources without sacrificing communicative effectiveness. This balance of clarity, structure, and brevity positions AutoForm as a powerful tool for facilitating intelligent agent communication in a variety of contexts.
@@ 6 Conclusion
In this work, we demonstrate that LLMs are capable of autonomously determining suitable non-NL formats for reasoning and communication, as facilitated by the AutoForm prompting method. Our analyses answer 5 key research questions, revealing that LLMs can generalize a reasoning format on a task from several task-specific examples and that the format is transferable across different LLMs. We also find that the communication formats generated by LLMs resemble traditional ACLs, offering both precision and efficiency. These insights advance our understanding of LLMs’ capabilities beyond NL, contributing to the enhancement of LLM reasoning and inter-agent communication.

### v2

@@ Abstract
Natural language (NL) has long been the predominant format for human cognition and communication, and by extension, has been similarly pivotal in the development and application of Large Language Models (LLMs). Yet, besides NL, LLMs have seen various non-NL formats during pre-training, such as code and logical expression. NL’s status as the optimal format for LLMs, particularly in single-LLM reasoning and multi-agent communication, has not been thoroughly examined. In this work, we challenge the default use of NL by exploring the utility of non-NL formats in these contexts. We show that allowing LLMs to autonomously select the most suitable format before reasoning or communicating leads to a 3.3 to 5.7% improvement in reasoning efficiency for different LLMs, and up to a 72.7% reduction in token usage in multi-agent communication, all while maintaining communicative effectiveness. Our comprehensive analysis further reveals that LLMs can devise a format from limited task instructions and that the devised format is effectively transferable across different LLMs. Intriguingly, the structured communication format decided by LLMs exhibits notable parallels with established agent communication languages, suggesting a natural evolution towards efficient, structured communication in agent communication. Our code will be released to facilitate further exploration.
@@
Weize Chen1thanks:   Equal Contribution, Chenfei Yuan [^1] footnotemark: 1, Jiarui Yuan [^1] footnotemark: 1, Yusheng Su1, Chen Qian1, Cheng Yang3thanks:   Corresponding author., Ruobing Xie2, Zhiyuan Liu122footnotemark: 2, Maosong Sun1 1 Tsinghua University 2 Tencent 3 Beijing University of Posts and Telecommunications {chenwz21,yuancf21,yuanjr22}@mails.tsinghua.edu.cn
@@ 1 Introduction
Recent research challenges the notion that NL is the ideal intermediate format for LLM reasoning and multi-agent communication. Emerging variants of CoT, such as Program-of-Thought (Chen et al., 2022; Gao et al., 2023) and X-of-Thought (Liu et al., 2023) have explored the use of alternative formats like code and mathematical equations, expanding the LLMs’ reasoning capabilities. However, these approaches often integrate external tools, where the alternative formats primarily serve as a means to facilitate tool execution (e.g., prompting LLM to generate code and use code interpreter execution result as the answer). This introduces complexity in discerning whether the performance improvements are attributable to the format itself or the accompanying tools. Additionally, while the natural ambiguities and emotions inherent in NL may align well with the nuances of human communication, these may not be desired in agent communication, where precision is more important. Nonetheless, current multi-agent research predominantly utilizes NL (Li et al., 2023a; Wu et al., 2023; Chen et al., 2023), with limited exploration of other potentially more accurate and efficient communication formats.
@@ 2 Related Work
Advances in Large Language Models (LLMs) have led to the development of autonomous agents like AutoGPT (Significant Gravitas, ) and OpenAI Assistant (OpenAI, 2023a), demonstrating success in diverse tasks (Shinn et al., 2023; Mialon et al., 2023; Zhou et al., 2023; Boiko et al., 2023). Recent research extends this to multi-agent systems for collaborative problem-solving (Du et al., 2023; Osika, 2023; Hong et al., 2023; Qian et al., 2023). CAMEL (Li et al., 2023a) explores collaborative problem-solving between two agents through role-playing. ChatEval (Chan et al., 2023) and PRD (Li et al., 2023b) assess model responses using multi-agent debates. AgentVerse (Chen et al., 2023) introduces a comprehensive framework for multi-agent collaboration, highlighting emergent inter-agent behaviors. However, the alternative formats of multi-agent communication remains underexplored and NL is directly adopted across various research. Pham et al. (2023) explores agent communication with hidden states, but is limited to agents with the same open-source LLM. Our work analyze communication formats for both homogeneous and heterogeneous LLMs.
@@ 4 Experiments
Model Logic Grid Coin Flip Info Essen MM QA AQuA Average
GPT-3.5 CoT 46.7±1.6 23.1±1.0 32.3±3.2 24.9±0.8 60.9±1.2 41.1±1.8
    +AutoForm 48.0±3.9 39.4±1.1 36.7±3.2 26.8±0.6 63.7±0.7 46.0±2.3
Gemini Pro CoT 49.7±0.2 47.5±0.2 34.3±0.7 28.1±0.7 56.3±0.6 43.2±0.5
    +AutoForm 51.2±0.8 57.6±0.7 39.2±1.8 31.3±1.1 60.0±0.4 47.9±1.1
GPT-4 CoT 61.8±1.6 93.4±1.0 78.4±2.5 38.4±1.1 79.1±0.3 71.8±1.5
    +AutoForm 65.8±2.2 98.4±0.2 76.9±2.5 41.7±0.9 80.4±0.8 74.1±1.6
@@ 4.1 Experimental Settings
Single-LLM Reasoning. In a preliminary experiment, we prompt various LLMs to use formats other than NL for solving reasoning problems. The results, shown in Table 6, reveal significant variability in performance across different formats. For instance, GPT-3.5 exhibited a 66.8% performance gap between using an ordered list and a multi-level list, highlighting the intrinsic suitability of specific formats for different tasks.
However, in practical applications, selecting optimal formats for each task may be impractical. To address this, we conduct a comprehensively evaluate the impact of automatically chosen thought formats on LLMs’ reasoning performance. We select reasoning benchmarks covering different types of reasoning, including logical reasoning (Logic Grid Puzzle (Srivastava et al., 2022) and Information Essentialy (Srivastava et al., 2022)), mathematical reasoning (AQuA (Ling et al., 2017)), causal reasoning (Minute Mysteries QA (Srivastava et al., 2022)) and symbolic reasoning (Coin Flip (Wei et al., 2022)). For all these tasks, we require the LLMs to generate the answer in a particular format, and we extract the answer with a written regular expression. The average accuracy and the standard deviation of each dataset over 3 runs are reported on most of the datasets. For more details on the experimental settings, please refer to Appendix A.
@@
Multi-Agent Communication. To measure whether alternative formats can streamline communication, we consider scenarios where two agents with different knowledge or contexts are tasked with answering a question. The answer to the question should be derived from the knowledge of only one of the agents, or both agents’ knowledge collectively, therefore requiring information exchange. The two agents speak in turn to discuss and give the final answer. To create such scenarios, we utilize three existing datasets: Hotpot QA (Yang et al., 2018), Wiki Hop (Welbl et al., 2018) and Narrative QA (Kociský et al., 2018). Hotpot QA and Wiki Hop are two multi-hop QA datasets, which require multiple sentences or paragraphs to deduce the final answer. We randomly assign the text segments provided in the datasets to two different agents. Communication is thus needed to derive the correct answer. We have also explored assigning each agent part of the supporting facts, thus needing more communication to derive the answer. Narrative QA requires the model to read the whole book and answer a question. The length of a book often exceeds the context limit of the LLMs. We split the books into nearly equal sizes for the two agents, and ask them to answer the question. For evaluation, we use RougeL as the primary metric. More details are elaborated in Appendix A.
To measure whether alternative formats can streamline communication, we consider scenarios where two agents with different knowledge or contexts are tasked with answering a question. The answer to the question should be derived from the knowledge of only one of the agents, or both agents’ knowledge collectively, therefore requiring information exchange. The two agents speak in turn to discuss and give the final answer. To create such scenarios, we utilize three existing datasets: Hotpot QA (Yang et al., 2018), Wiki Hop (Welbl et al., 2018), and Narrative QA (Kociský et al., 2018). Hotpot QA and Wiki Hop are two multi-hop QA datasets, which require multiple sentences or paragraphs to deduce the final answer. We randomly assign the text segments provided in the datasets to two different agents. Communication is thus needed to derive the correct answer. We also explore assigning each agent part of the supporting facts, thus needing more communication to derive the answer. Narrative QA requires the model to read the whole book and answer a question. The length of a book often exceeds the context limit of the LLMs. We split the books into nearly equal sizes for the two agents and ask them to answer the question. While supporting facts in Narrative QA are not guaranteed to reside in different segments, this division still introduces critical challenges: agents must identify who possesses the necessary information, necessitating communication. For evaluation, we use RougeL as the primary metric. More details are elaborated in Appendix A.
Research Questions. To comprehensively explore the potential of LLMs in selecting suitable formats for both single-agent reasoning and multi-agent communication, we conduct an in-depth analysis addressing six pivotal research questions (RQs). These questions aim to unravel the intricacies of format selection by LLMs and its impact on task performance across various scenarios:
  • RQ1: Can LLMs select the suitable formats autonomously? (Section 4.2)
  • RQ2: What formats are chosen in single-LLM reasoning? (Section 4.3)
  • RQ3: Can LLMs devise a general format for a task based on some task inputs? (Section 4.4)
  • RQ4: Is the decided format transferable between different LLMs? (Section 4.5)
  • RQ5: What are the features of the formats used in multi-agent communication?  (Section 4.6)
  • RQ6: Does the autonomously determined multi-agent communication format align with conventional agent communication languages such as KQML (Finin et al., 1994)? (Section 4.7)
4.2 RQ1: The Capability of LLMs on Selecting Suitable Format
Single-LLM Reasoning. The comparative efficacy of the AutoForm approach over the conventional Chain-of-Thought (CoT) methodology in single-LLM reasoning tasks is encapsulated in Table 1. We observe clear performance improvements when employing AutoForm across different datasets, compared to the baseline CoT method.
@@ 4.2 RQ1: The Capability of LLMs on Selecting Suitable Format
Multi-Agent Communication. The outcomes of our multi-agent communication experiments, detailed in Table 2, provide valuable insights into the efficiency and effectiveness of utilizing alternative communication formats in collaborative environments. In this experiment, we experiment with different model pairings to explore this robustness more comprehensively. Since the initial speaking agent often sets the tone for the communication format, we vary the speaking order in pairings of heterogeneous models, e.g., GPT-4 + GPT-3.5 in the table indicates GPT-4 is the initiator. Due to page limit, we place results where GPT-3.5 initiates the conversation, and where the supporting facts are split and distributed to two agents at Appendix C.
@@
4.3 RQ2: Formats Chosen in Single-LLM Reasoning
@@
Model Logic Grid Coin Flip Info Essen MM QA AQuA Average
GPT-3.5 (Instance-Based) 48.0 49.2 39.7 26.1 66.5 45.9 (+7.5%)
GPT-3.5 (Task-Based) 51.0 62.8 42.6 28.1 65.0 49.9 (+11.5%)
Gemini Pro (Instance-Based) 39.5 44.8 35.3 27.1 59.4 41.2 (+2.5%)
Gemini Pro (Task-Based) 41.5 47.8 35.3 28.6 59.4 42.5 (+3.8%)
GPT-4 (Instance-Based) 71.5 100.0 76.5 41.4 78.3 73.5 (+3.2%)
GPT-4 (Task-Based) 70.0 99.8 75.0 41.4 79.5 73.1 (+2.8%)
GPT-4 $\Rightarrow$ GPT-3.5 (Task-Based) 47.5 83.0 35.3 23.2 59.4 49.7 (+11.2%)
GPT-4 $\Rightarrow$ Gemini Pro (Task-Based) 40.0 74.6 25.0 25.6 47.2 42.5 (+3.8%)
GPT-3.5 $\Rightarrow$ GPT-4 (Task-Based) 65.5 98.6 72.1 42.9 79.5 71.7 (+1.3%)
4.4 RQ3: Generalization of Format Selection Based on Task Inputs
@@
4.5 RQ4: Transferability of Format Across Different LLMs
@@
The results are presented in the last group of Table 3. Generally, the format is transferable, but may lead to slightly inferior performances compared to the homogeneous setting. For example, transferring the format decided by GPT-4 to GPT-3.5 or Gemini Pro leads to a decrease compared to the homogeneous setting for the two LLMs in most of the tasks except for Coin Flip. On the other hand, when transferring the format decided by GPT-3.5 to GPT-4, the results are generally comparable to the homogeneous setting for GPT-4. For information essentiality, the format selected by GPT-3.5 is generally less efficient, leading to incomplete problem-solving processes. We give examples in Tables 33, 39 and 36. For other tasks, the format generated by GPT-3.5 proves adequate and is easily interpreted by GPT-4, resulting in similar performances to the homogeneous setting for GPT-4.
4.6 RQ5: Features of Communication Format
@@
Brevity and Efficiency. Another key feature is the focus on brevity, enhancing communication efficiency. Formats chosen by LLMs often omit elements like pleasantries or emotive expressions, resulting in concise exchanges. This brevity conserves computational resources and concentrates dialogue on the task at hand, optimizing the communication process for faster, more efficient information exchange. This is particularly beneficial in scenarios requiring rapid and effective decision-making.
4.7 RQ6: Alignment with Conventional Agent Communication Languages
@@
Hotpot QA Wiki Hop Narrative QA
@@
These results highlight two key implications. First, AutoForm can generate communication patterns similar to established ACLs. Second, it efficiently distills these traditional formats into a more concise form, conserving computational resources while maintaining communicative effectiveness. This balance of clarity, structure, and brevity makes AutoForm a powerful tool for facilitating intelligent agent communication in various contexts.
@@ 5 Conclusion
In this work, we demonstrate that LLMs can autonomously determine suitable non-NL formats for reasoning and communication using the AutoForm prompting method. Our analyses address six key research questions, showing that LLMs can generalize a reasoning format from task-specific examples and transfer it across different models. Additionally, the communication formats generated by LLMs resemble traditional ACLs, offering both precision and efficiency. These insights enhance our understanding of LLMs’ capabilities beyond NL, improving LLM reasoning and inter-agent communication.